# FreeDiff: Progressive Frequency Truncation for Image Editing with Diffusion Models

### v1

@@ Appendix 0.A More intermediate features from generation process
Appendix 0.B Editing difference from the frequency perspective
Examples of various editing types applied to different images using two attention-based methods, P2P[6] and PNP[21], and direct editing are shown in Fig. 9. These examples support our hypothesis that direct editing inadvertently introduces an excess of low-frequency components, due to the learned prior and weighting schedule of the denoising network, leading to an undesired alteration in non-target regions.
Figure 9: Editing results from attention-based refining methods P2P[6]+NTI[12], PNP[21]+fixed-point inversion and directly applying guidance. Column d) and e) shows $\mathcal{F}_{diff}(I_{src},I_{edit})$ between <source image, attention-based editing>, <source image, direct editing>, respectively. The $\mathcal{F}_{diff}(I_{src},I_{edit})$ is normalized to the same numerical scale in each row.
Appendix 0.C Qualitative and quantitative results
@@ 0.C.1 PIE Dataset
These two problems hinder the accuracy of evaluation, since for attention-based methods, the best default hyperparameter sets vary largely for different editing types. For our method, accurately defining the editing types is crucial for selecting appropriate reference hyperparamters. Consequently, we will re-categorize the PIE dataset in the future, which will be detailed in the next section.
0.C.2 Editing categories and hyperparameters with FreeDiff
@@
0.C.3 Two-step process for editing colors and environments
@@
0.C.4 Quantitative results
As our method largely relies on distinct hyperparameter sets for precise editing refinement, evaluating its efficacy under the original PIE categories does not fully showcase its potential. In light of the challenges within the PIE datasets and the characteristics of FreeDiff, we have conducted our quantitative analysis on selected categories of the PIE dataset, including objects changing, adding, deleting, and material alterations. We selected the entire category, excluding any misclassified images and instances where fixed-point inversion fails to reconstruct the original image. Additionally, we chose the former image-text pairs from that category. In total, 208 images were selected.
Table 2: Quantitative results from partial dataset of PIE[10]
Ours P2P PNP MasaCtrl
CLIP Score (whole image) $\uparrow$ 25.5140 24.7521 25.4717 24.6614
Background LPIPS $\downarrow$ 11.14 11.83 15.01 13.97
We evaluated the CLIP score across the entire image and the LPIPS score for the background region, with results detailed in Tab. 2. While our method exhibits slightly better over other attention-based editing methods, we do not consider the CLIP score to be an effective metric for evaluating editing quality. This is because, according to human perception, the editing results produced by P2P are generally better than those by PNP.
0.C.5 Qualitative results
@@
Figure 10: Qualitative comparisons in changing materials, altering poses, and shapes using images from the PIE dataset [10]. The analysis juxtaposes our approach with 3 typical attention-based editing methods: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.
Figure 11: Qualitative comparisons in changing styles, colors, and environment using images from the PIE dataset [10]. The analysis juxtaposes our approach with 3 typical attention-based editing methods: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.
Figure 12: Qualitative comparisons in removing objects using images from the PIE dataset [10]. The analysis juxtaposes our approach with 3 typical attention-based editing methods: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.
Figure 13: Qualitative comparisons in adding objects using images from the PIE dataset [10]. The analysis juxtaposes our approach with 3 typical attention-based editing methods: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.
Figure 14: Qualitative comparisons in changing identities, altering shape, and adding objects using images from the PIE dataset [10]. The analysis juxtaposes our approach with 3 typical attention-based editing methods: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.

### v2

@@ Appendix 0.B More intermediate features from generation process
Appendix 0.C Editing difference from the frequency perspective
Examples of various editing types applied to different images using two ABMs, P2P[6] and PNP[21] , and direct editing are shown in Fig. 9. These examples support our hypothesis that direct editing inadvertently introduces an excess of low-frequency components, due to the learned prior and weighting schedule of the denoising network, leading to an undesired alteration in non-target regions.
Figure 9: Editing results from ABMs: P2P [6] +NTI[12] , PNP[21] +fixed-point inversion and directly applying guidance. Column d) and e) shows $\mathcal{F}_{diff}(I_{src},I_{edit})$ between <source image, attention-based editing>, <source image, direct editing>, respectively. The $\mathcal{F}_{diff}(I_{src},I_{edit})$ is normalized to the same numerical scale in each row.
Appendix 0.D Qualitative and quantitative results
@@ 0.D.1 PIE Dataset
These two problems hinder the accuracy of evaluation, since for ABMs, the best default hyperparameter sets vary largely for different editing types. For our method, accurately defining the editing types is crucial for selecting appropriate reference hyperparamters. Consequently, we will re-categorize the PIE dataset in the future, which will be detailed in the next section.
0.D.2 Editing categories and hyperparameters with FreeDiff
@@
0.D.3 Two-step process for editing colors and environments
@@
0.D.4 Quantitative results
For the overall quantitative results on the partial PIE dataset shown in Tab.1, we selected editing types where the comparison attention-based methods(ABMs) tended to perform well (change, add, and delete objects, change materials and poses). We did not include image where the inversion failed, or the ABMs catastrophically failed. Additionally, for most categories, we chose the former half of image-text pairs for the partial dataset. We did not cherry pick the images to improve our method’s results. When testing the ABMs, we found some methods had a high number of failure cases on the claimed specialized type. Finally, we did not include editing types (style and color change) where ABMs required a large search to fine-tune the hyperparameters, since this is infeasible due to some method’s computational complexity and lack of guidelines for searching, if we want to compare the results with our best hyperparameters. In total, 208 images were selected.
For the sub-categories results shown in Tab.2, we further correct the partial dataset from the mentioned issues, and 203 images are selected.
Overall, our experiments are conducted fairly since we select the images that the comparison ABMs perform well on. All editing results and hyperparams will be released with the source code.
We evaluated the CLIP score across the entire image and the LPIPS score for the background region, with results detailed in Tab. 1. While our method exhibits slightly better over other ABMs, we do not consider the CLIP score to be an effective metric for evaluating editing quality. This is because, according to human perception, the editing results produced by P2P are generally better than those by PNP.
0.D.5 Qualitative results
@@
Figure 10: Qualitative comparisons in changing materials, altering poses, and shapes using images from the PIE dataset[10] . The analysis juxtaposes our approach with 3 typical ABMs: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.
Figure 11: Qualitative comparisons in changing styles, colors, and environment using images from the PIE dataset[10] . The analysis juxtaposes our approach with 3 typical ABMs: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.
Figure 12: Qualitative comparisons in removing objects using images from the PIE dataset[10] . The analysis juxtaposes our approach with 3 typical ABMs: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.
Figure 13: Qualitative comparisons in adding objects using images from the PIE dataset[10] . The analysis juxtaposes our approach with 3 typical ABMs: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.
Figure 14: Qualitative comparisons in changing identities, altering shape, and adding objects using images from the PIE dataset[10] . The analysis juxtaposes our approach with 3 typical ABMs: P2P, PNP, and MasaCtrl. Direct editing results with fixed-point inversion are also included as a baseline for benchmarking.